A ACETIC ACID SURVEY TOWARD HIGH- AND LOW-MASS STAR FORMING REGIONS

\underline{Y.-S. SHIAO}, L. W. LOONEY, L. E. SNYDER, D. N. FRIEDEL, Department of Astronomy, University of Illinois at Urbana-Champaign, IL 61801; A. J. REMIJAN, National Radio Astronomy Observatory, Charlottesville, VA 22903.

We have used the Combined Array for Research in Millimeter-wave Astronomy (CARMA) to conduct a survey of acetic acid (CH_{3}COOH) at 3 mm wavelengths toward the hot core G19.61-0.23 and the two deeply embedded low-mass protostars IRAS 16293-2422 and NGC 1333 IRAS 4. The observed CH_{3}COOH transitions are the 10_{∗,10} − 9_{∗,9} E and A lines. We have detected (>3 σ level) these two lines in G19.61-0.23. The abundance ratio of CH_{3}COOH to its isomer, methyl formate (HCOOCH_{3}), is comparable to other CH_{3}COOH sources. However, the CH_{3}COOH lines in IRAS 16293-2422 and NGC 1333 IRAS 4 are not detected in our observations. We set upper limits for the column densities for each source. Based on the CH_{3}COOH detection in IRAS 16293-2422 reported by Cazaux et al., we should have detected the two lines. Therefore, we suggest that the CH_{3}COOH detection in IRAS 16293-2422 is still uncertain.

We acknowledge support from the Laboratory for Astronomical Imaging at the University of Illinois and NSF grant AST-0540459.